# Orbital Dynamics and Gravitational Wave Signatures of Extreme Mass Ratio Inspirals in Galactic Dark Matter Halos

### 02 Mar 2026 v1

@@ I Introduction
The continued success of ground-based gravitational wave (GW) observatories has established the era of GW astronomy [1, 2, 3, 4, 5, 6]. To extend the observational window into the millihertz frequency band [7, 8, 9, 10], space-based gravitational wave missions—represented by the Laser Interferometer Space Antenna (LISA) [11, 8], Taiji [12, 13], and TianQin [14, 15]—are currently under active development. As one of the most promising sources for these space-based detectors, extreme mass ratio inspirals (EMRIs) are typically characterized by a stellar-mass compact object (with a mass $m$ ranging from $1M_{\odot}$ to $100M_{\odot}$ approximately), orbiting a supermassive black hole (SMBH) with a mass $M$ spanning from $10^{5}M_{\odot}$ to $10^{8}M_{\odot}$, which generally resides in galactic nuclei [16]. Given this extreme mass ratio, the stellar-mass compact object can effectively behave as a test particle, tracing out approximately $10^{4}\sim 10^{5}$ orbital cycles within the detector’s sensitive frequency band. Crucially, these very high precision observations encode rich information, enabling us to decode the astrophysical environments (such as accretion disks, dark matter, and electromagnetic fields) of the black hole from GW signatures [17, 18, 19, 20, 21, 22]. By analyzing the GW signals of EMRIs, it is possible to directly constrain the influences of these environments.
@@
A vast number of astrophysical observations, including the cosmic microwave background [23], gravitational lensing and galaxy rotation curves [24, 25], consistently indicate that the universe is dominated by dark matter and dark energy. In galactic nuclei, dark matter is expected to form halo structures around the central SMBH, implying that these SMBHs are embedded in dense matter environments [26, 27]. Extensive numerical simulations and observational studies have motivated various theoretical models to describe these dark matter density profiles, including the Navarro-Frenk-White (NFW) [28, 29], Beta  [30, 31], Moore [32], Einasto [33, 34], Dehnen [35], Hernquist [36] and Burkert [37] models. The presence of such a halo inevitably influences the orbital dynamics of EMRIs through two mechanisms: firstly, the gravitational potential contribute from the dark matter halo modifies the background spacetime geometry; secondly, interactions between the compact object and the surrounding medium give rise to dissipative effects, most notably dynamical friction [38, 39] and accretion [40, 41, 42, 43]. Over the long-duration inspiral evolution, these environmental effects continuously reduce the orbital energy and angular momentum, causing a cumulative “dephasing” of the gravitational waveform relative to vacuum predictions. This distinctive phase shift provides a critical observational signature for constraining the dark matter halo models and revealing the matter distribution in galaxies [44, 45, 46, 47, 48]. Although the influence of dark matter environments on EMRI dynamics has been investigated [49, 50, 45, 51, 52, 53, 54], distinguishing between different halo density profiles through GW observations remains a significant theoretical challenge.
@@
The dark matter effects on EMRI systems and their GW signatures can be extracted through the investigations of precessing Keplerian orbits [55, 48, 56, 57], periodic orbits (with zoom-whirl-vertex behaviors) [58, 59, 60], and quasi-normal modes [44, 61, 62, 63]. In our previous study, the periodic orbits around an SMBH influenced by various dark matter halo models were investigated for fixed orbital energies or angular momenta (without considering any dissipation mechanisms) [58]. The periodic orbital shapes and GW waveforms predicted by NFW and Beta halo models exhibit nearly identical characteristics for certain dark matter parameters, making them observationally indistinguishable in such non-dissipative assumptions. Consequently, analyses limited to conservative dynamics or short-term evolution are insufficient to identify the specific dark matter distribution from GW observations. To resolve this degeneracy, it is necessary to explore the secular orbital evolution of the EMRI system over a large number of orbital periods with dissipative mechanisms taken into account, i.e. gravitational radiation backreaction, dynamical friction. By monitoring the cumulative phase shift induced by gravitational radiation, dynamical friction, and accretion over a long-duration inspiral, it becomes possible to distinguish these halo models.
@@
In this work, we aim to resolve the aforementioned NFW/Beta degeneracy by modeling the long-term adiabatic evolution of EMRIs affected by dark matter environments (described by NFW and Beta density profiles) [64, 65, 66, 67, 68, 69]. Specifically, we focus on quantifying how gravitational radiation reaction, accretion, and dark matter dynamical friction induce the variation of orbital energies, angular momenta, eccentricity, semi-latus rectum, and orbital phase shifts, enabling observational differentiation between these density profiles. We employ the Numerical Kludge method to generate waveforms from these evolved trajectories and calculate the cumulative dephasing over a typical one-year observation timescale [70], from which a systematic comparison on phase evolution in NFW and Beta halo models can be carried out to magnify their distinguishability. The primary goal is to determine the conditions under which space-based detectors can resolve the specific density structure of the galactic dark matter halo.
@@ II.1 Spacetime Metric
Considering the dark matter halo as an extended mass distribution surrounding the black hole, the total gravitational field can be treated as a combination of the SMBH’s gravitational field and the dark matter halo’s field. Consequently, the total metric function $f(r)$ is decomposed into the contribution from the central black hole and a correction term $f_{DM}(r)$ arising from the halo [71] [^1] It can be verified that the spacetime metric constructed via the decomposition in Eqs. (2) and (4) satisfies the Einstein field equations, with the $T_{00}$ component of the energy-momentum tensor specified by the dark matter density profile [71, 72, 73]:
$f(r)=f_{DM}(r)-\frac{2M}{r}.$ (2)
@@
To obtain the explicit metric forms, specific dark matter density profiles $\rho(r)$ are required. We adopt two widely used profiles: the Navarro-Frenk-White (NFW) model and the Beta model. The density distribution functions for these models are defined as [28, 29, 31]:
@@ IV Gravitational Waveforms and Phase Accumulation
Having assessed the long-term orbital evolution under different dark matter scenarios, we now focus on the corresponding gravitational wave (GW) signatures. This section presents a comparative analysis of the long-term GW signals from EMRIs within dark matter environments, concentrating on the waveform modulations and phase accumulations affected by the Beta and NFW models. To simulate the GW signals, we employ the Numerical Kludge method [70]. In the quadrupole approximation, the spatial metric perturbation $h_{ij}$ is determined by the second time derivative of the source’s mass quadrupole moment. For a compact object of mass $\mu$ orbiting a central black hole of mass $M$, this perturbation can be expressed in terms of the orbital kinematics as [91]:
@@
where $D_{L}$ represents the luminosity distance from the EMRI source to the detector, $v_{i}$ is the velocity vector of the SCO, $r$ represents the radial distance of the SCO from the central SMBH, and $n_{i}=x_{i}/r$ is the unit direction vector. To characterize the GW signal as perceived by the detector, we establish a detector-adapted coordinate system $(X,Y,Z)$. The transformation from the source frame is parameterized by the inclination angle $\iota$ and the pericenter longitude $\zeta$. Explicitly, the orthonormal basis vectors for the detector frame are constructed as follows [91, 92]:
@@
In this coordinate system, the gravitational wave polarizations $h_{+}$ and $h_{\times}$ observed at the detector are defined as:

### 12 Mar 2026 v2

@@ I Introduction
The continued success of ground-based gravitational wave (GW) observatories has established the era of GW astronomy [1, 2, 3, 4, 5, 6]. To extend the observational window into the millihertz frequency band [7, 8, 9, 10], space-based gravitational wave missions—represented by the Laser Interferometer Space Antenna (LISA) [11, 8], Taiji [12, 13], and TianQin [14, 15]—are currently under active development. As one of the most promising sources for these space-based detectors, extreme mass ratio inspirals (EMRIs) are typically characterized by a stellar-mass compact object (with a mass $m$ ranging from $1M_{\odot}$ to $100M_{\odot}$ approximately), orbiting a supermassive black hole (SMBH) with a mass $M$ spanning from $10^{5}M_{\odot}$ to $10^{8}M_{\odot}$, which generally resides in galactic nuclei [16]. Given this extreme mass ratio, the stellar-mass compact object can effectively behave as a test particle, tracing out approximately $10^{4}\sim 10^{5}$ orbital cycles within the detector’s sensitive frequency band. Crucially, the high precision observations on EMRIs encode rich information, enabling us to decode the astrophysical environments (such as accretion disks, dark matter, and electromagnetic fields) near the SMBH from GW signatures [17, 18, 19, 20, 21, 22, 23, 24]. By analyzing the GW signals, it is possible to directly constrain the influences of these environments. Additionally, EMRIs present exceptional value in probing the quantum gravity effects beyond general relativity [25, 26, 27, 28].
@@
A vast number of astrophysical observations, including the cosmic microwave background [29], gravitational lensing and galaxy rotation curves [30, 31], consistently indicate that the universe is dominated by dark matter and dark energy. In galactic nuclei, dark matter is expected to form halo structures around the central SMBH, implying that these SMBHs are embedded in dense matter environments [32, 33]. Extensive numerical simulations and observational studies have motivated various theoretical models to describe these galactic dark matter halos, including the Navarro-Frenk-White (NFW) [34, 35, 36], Beta [37], Moore [38], Einasto [39, 40], Dehnen [41], Hernquist [42] and Burkert [43] models. The presence of dark matter halos inevitably exerts profound impacts on gravitational systems in galaxies, such as binary black hole (BBH) coalescences, intermediate mass ratio inspirals (IMRIs) and extreme mass ratio inspirals (EMRIs) [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Particularly, the dark matter halo influences the orbital dynamics of EMRIs (or IMRIs) through two mechanisms: firstly, the gravitational potential contribute from the dark matter halo modifies the background spacetime geometry; secondly, interactions between the compact object and the surrounding medium give rise to dissipative effects, most notably dynamical friction [55, 56] and accretion [57, 58, 59, 60]. Over the long-duration inspiral evolution, these environmental effects continuously reduce the orbital energy and angular momentum, causing a cumulative “dephasing” of the gravitational waveform relative to vacuum predictions. This distinctive phase shift provides a critical observational signature for constraining the dark matter halo models and revealing the matter distribution in galaxies [61, 62, 63, 64].
@@
The dark matter effects on EMRI systems and their GW signatures can be extracted through the investigations of precessing Keplerian orbits [65, 66], periodic orbits (with zoom-whirl-vertex behaviors) [67, 68, 69], chaotic dynamics [70, 71], and quasi-normal modes [72, 73, 74, 75]. Although the influence of dark matter environments on EMRI dynamics has been investigated, distinguishing between different halo density profiles through GW observations remains a significant theoretical challenge. In our previous study, the periodic orbits around an SMBH influenced by various dark matter halo models were investigated for fixed orbital energies or angular momenta (without considering any dissipation mechanisms) [67]. The periodic orbital shapes and GW waveforms predicted by NFW and Beta halo models exhibit nearly identical characteristics for certain dark matter parameters, making them observationally indistinguishable in such non-dissipative assumptions. Consequently, analyses limited to conservative dynamics or short-term evolution are insufficient to identify the specific dark matter distribution from GW observations. To resolve this degeneracy, it is necessary to explore the secular orbital evolution of the EMRI system over a large number of orbital periods with dissipative mechanisms taken into account, i.e. gravitational radiation backreaction, dynamical friction. By monitoring the cumulative phase shift induced by gravitational radiation, dynamical friction, and accretion over a long-duration inspiral, it becomes possible to distinguish these halo models.
@@
In this work, we aim to resolve the aforementioned NFW/Beta degeneracy by modeling the long-term adiabatic evolution of EMRIs affected by dark matter environments (described by NFW and Beta density profiles). Specifically, we focus on quantifying how gravitational radiation reaction, accretion, and dark matter dynamical friction induce the variation of orbital energies, angular momenta, eccentricity, semi-latus rectum, and orbital phase shifts, enabling observational differentiation between these density profiles. We employ the adiabatic approximation [76, 77, 78, 79, 80, 81] to evaluate the long-term orbital evolution and use the Numerical Kludge method [82, 83, 84, 85] to generate waveforms from these evolved trajectories and calculate the cumulative dephasing over a typical one-year observation timescale, from which a systematic comparison on phase evolution in NFW and Beta halo models can be carried out to magnify their distinguishability. The primary goal is to determine the conditions under which space-based detectors can resolve the specific density structure of the galactic dark matter halo.
@@ II.1 Spacetime Metric
Considering the dark matter halo as an extended mass distribution surrounding the black hole, the total gravitational field can be treated as a combination of the SMBH’s gravitational field and the dark matter halo’s field. Consequently, the total metric function $f(r)$ is decomposed into the contribution from the central black hole and a correction term $f_{\text{DM}}(r)$ arising from the halo [86] [^1] It can be verified that the spacetime metric constructed via the decomposition in Eqs. (2) and (4) satisfies the Einstein field equations, with the $T_{00}$ component of the energy-momentum tensor specified by the dark matter density profile [86, 87, 88]:
$f(r)=f_{\text{DM}}(r)-\frac{2M}{r}.$ (2)
@@
To obtain the explicit metric forms, specific dark matter density profiles $\rho(r)$ are required. We adopt two widely used profiles: the Navarro-Frenk-White (NFW) model and the Beta model. The density distribution functions for these models are defined as [34, 35, 36, 37]:
@@ IV Gravitational Waveforms and Phase Accumulation
Having assessed the long-term orbital evolution under different dark matter scenarios, we now focus on the corresponding gravitational wave (GW) signatures. This section presents a comparative analysis of the long-term GW signals from EMRIs within dark matter environments, concentrating on the waveform modulations and phase accumulations affected by the Beta and NFW models. To simulate the GW signals, we employ the Numerical Kludge method [84, 85]. The main strategy of the Kludge method is as follows: first, calculate the orbit of the SCO by solving the geodesic equation, and then use the quadrupole formula of gravitational radiation to get the corresponding gravitational waves[84, 105]. In the quadrupole approximation, the spatial metric perturbation $h_{ij}$ is determined by the second time derivative of the source’s mass quadrupole moment. For a compact object of mass $\mu$ orbiting a central black hole of mass $M$, this perturbation can be expressed in terms of the orbital kinematics as [106, 107, 108]:
@@
where $D_{L}$ represents the luminosity distance from the EMRI source to the detector, $v_{i}$ is the velocity vector of the SCO, $r$ represents the radial distance of the SCO from the central SMBH, and $n_{i}=x_{i}/r$ is the unit direction vector. To characterize the GW signal as perceived by the detector, we establish a detector-adapted coordinate system $(X,Y,Z)$. The transformation from the source frame is parameterized by the inclination angle $\iota$ and the pericenter longitude $\zeta$. Explicitly, the orthonormal basis vectors for the detector frame are constructed as follows:
@@
In this coordinate system, the gravitational wave polarizations $h_{+}$ and $h_{\times}$ observed at the detector are expressed as [107, 108]:
@@ V Conclusion
Acknowledgements.
This research is supported by the Natural Science Foundation of China (Grant No. 12175212), Natural Science Foundation of Chongqing Municipality (Grant No. CSTB2022NSCQ-MSX0932), the Scientific and Technological Research Program of Chongqing Municipal Education Commission (Grant No. KJQN202201126).